Things did not go quite as planned. You invested $10,250.00, part of it in stock that paid 12.5% annual interest. However, the rest of the money suffered a 6.7% loss. If the total annual income from both investments was $346.00, how much was invested at each rate?
Answer by stanbon(75887)

You invested $10,250.00, part of it in stock that paid 12.5% annual interest. However, the rest of the money suffered a 6.7% loss. If the total annual income from both investments was $346.00, how much was invested at each rate?
----
Quantity Eq::: x + y = 10,250
Interest Eq::0.125x - 0.067y = 346
-----------------------------------
Multiply thru the Interest equation by 1000
Multiply thru the Quantity Eq. by 125
--------------------------
125x + 125y = 125*10,250
125x - 67y = 346000
---------------------------------------
Subtract and solve for "y":
192y = $935,250
y = $4871.09 (amt. invested at 6.7%
Solve for "x":
x + y = 10,250
x + 4671.09 = 10,250
x = $5378.91 (amt. invested at 12.5%
=========================================
Cheers,
Stan H.
==============